princeevee:

Suppose the test scores of students in a class are normally distributed with a mean of 92 and a standard deviation of 3. What is the z-score for a student who scored 86 on a test?

Ultrilliam:

@Vocaloid

Vocaloid:

z-score for a particular value: (that value - mean)/standard deviation

Vocaloid:

(86-92)/(3) = ?

princeevee:

-2. thanks

princeevee:

Suppose the velocities of golf swings for amateur golfers are normally distributed with a mean of 95 mph and a standard deviation of 3.9 mph. What is the difference in velocities between a golfer whose z-score is 0 and another golfer whose z-score is −1?

Vocaloid:

work backwards using the z-score formula to find the values

Vocaloid:

(x - 95)/3.9 = 0 (y - 95)/3.9 = -1 calculate x and y

princeevee:

um....

Vocaloid:

(x - 95)/3.9 = 0 multiply both sides by 3.9, what do you get?

princeevee:

i got it actually

Vocaloid:

ok great